Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

night-report-page-20131112

Observers: Susanna Spiro        TO: Monica Castillo
Support team: S.Benetti, N.Elias-Rosa
TAT on duty: Joe Anderson, Matt Nicholl
------------------------------------------------

Subrun 3, night 3

Instruments: SOFI, EFOSC2

WEATHER CONDITION
beginning: wind speed 6.7 m/s; humidity 10 % ; clear
middle: wind speed  4.6 m/s; humidity 19 %;  clear
end: wind speed  2 m/s; humidity 16 %;  clear

afternoon CALIBRATIONS

PESSTO_SOFI_FlatJHK 
PESSTO_SOFI_specflat
PESSTO_SOFI_SpecArc

OBSERVATIONS: started delayed of half an hour due to technical problems with the guide star
Target                       Type             UT          Filter/Grism      Slit            Seeing    AIrmass    Grade (A/B/C)     Comments

Hip094570              tellSTD         00:30           BG,RG           1.0             0.7          1.55              A             
SN2013fq                follow-up     00:45           JHKs                               0.6           1.5               A            
SN2013fq                follow-up     01:00           BG+RG         1.0             0.6           1.6               A
Hip110035              tellSTD         02:10           BG,RG           1.0             0.8           1.5               A
SN2013fs                follow-up     02:15           JHKs                                0.7           1.5               A
SN2013fs                follow-up     02:45           BG+RG          1.0             0.7           1.7               A        RG stopped  4/8
Hip000682              tellSTD         04:35           BG,RG           1.0             0.8           1.8               A
PSNJ01591443+1901339  follow-up       04:45        JHKs                        0.9           1.6               ?         problem in the coadded images in H&Ks, reason unknown, probably due to an error of the pipeline?  
SN2013ej                follow-up     05:53            JHKs                              0.6            2.2               A
SN2013fc                follow-up     06:02            JHKs                              0.6            1.2               A
LTT3218                 fluxSTD        07:08          BG,RG            1.0             0.7            1.2               A    

EFOSC2
LTT3864                 STD              08:00             gr13              1.0             0.7            1.5            A
SN2013fc                follow-up     08:10           UBVRi                                0.8            1.6            A
LSQ13cve                classification  08.36           gr13              1.0             0.6            1.7            A     type Ia 

skyflat grz 

After changing from SOFI to EFOSC2, there were some communication problem with P2PP: the Night Report from the TO does not contain the informations about the observations with EFOSC2. 
Comments